# T versus CP effects in DUNE and T2HK

### v1

@@ 3 T violation from $L$ dependence
Further details on bin-wise and individual sensitivities are given in appendix A. Let us emphasize the crucial importance of the low-energy tail of the DUNE event spectrum for this analysis, covering the 2nd oscillation maximum. It is well known that the second oscillation maximum provides high sensitivity to $\delta_{\rm CP}$, see, for instance,[37, 34, 33, 32, 30, 28]. Here we argue, that this is a manifestation of T violation, as no neutrino/anti-neutrino comparison is involved but only the $L$ dependence of the oscillation probability (see further discussions in section 4 and appendix A).
Figure 1 illustrates also the dramatic impact of prior knowledge of the oscillation parameters on the sensitivity. For the green curves, no prior information on the mixing angles is assumed and we obtain a marginal sensitivity well below $2\sigma$. This situation is similar (though not identical) to the model-independent approach of our previous paper [29]. [^1] Note that in [29] we have assumed a DUNE exposure of 840 kt MW yr in order to boost the sensitivity, compared to the default exposure of 336 kt MW yr assumed here. In contrast, if a prior corresponding to the uncertainties from the current global fit on $\sin^{2}\theta_{ij}$ is imposed (blue curves in fig. 1), the sensitivity increases drastically, reaching more than $3\sigma$ for $\delta_{\rm CP}\approx 90^{\circ}$ for individual bins and more than $4\sigma$ for the combination. We note that at the time of the DUNE and T2HK experiments, an even better accuracy would be available from JUNO on $\theta_{12}$ and $\theta_{13}$ [47] and from the DUNE/T2HK disappearance channels on $\theta_{23}$. Comparison of the blue and red curves in fig. 1 shows that a perfect knowledge of the mixing angles leads only to a slight improvement compared to present uncertainties.
@@
The left panel of fig. 3 shows that indeed the knowledge of all three mixing angles is important to reach such good sensitivities as displayed in fig. 1, especially on $\theta_{13}$ and $\theta_{23}$. Furthermore, an important effect comes also from the $\theta_{23}$ octant degeneracy. In our default analysis we assume a best fit value in the second octant, $\sin^{2}\theta_{23}=0.561$, and impose a prior on $\sin^{2}\theta_{23}$, which does determine the octant and therefore breaks the well-known octant degeneracy [50]. In the right panel of fig. 3 we study this issue in some detail. For the red curves we assume a prior on $\sin^{2}\theta$ (breaking the degeneracy), whereas for the blue curves the prior is imposed on $\sin^{2}2\theta$, which is symmetric with respect to the octant. By comparing the red and blue curves we see that the knowledge of the $\theta_{23}$ octant is rather important, in particular for the two combinations of true values ($\theta_{23}>45^{\circ}$, $\delta_{\rm CP}\simeq 90^{\circ}$) and ($\theta_{23}<45^{\circ}$, $\delta_{\rm CP}\simeq 270^{\circ}$). If the octant is known, the sensitivity is somewhat better for $\theta_{23}^{\rm true}<45^{\circ}$. While fig. 3 emphasizes the important to resolve the octant degeneracy, we note that the experiments considered here have good sensitivity to indeed lift this degeneracy, see e.g. [51, 52, 12, 53]. Therefore, we adopt the prior on $\sin^{2}\theta_{23}$ as our default assumption, assuming that the octant is known by the time the data becomes available.
@@
As the considered energy range is in the low-energy tail for DUNE, small event numbers are expected. Therefore, we study the sensitivity as a function of the DUNE exposure in fig. 4, together with the assumed energy resolution. We have checked that the results depend only mildly on both, the exposure and energy resolution in T2HK. For our default assumption on the energy resolution [42] as quoted in section 2.2 (blue curve), sensitivities above $4\sigma$ are obtained for exposures above 250 (400) kt MW yr in the neutrino running mode for a true $\delta_{\rm CP}=90^{\circ}$ ($270^{\circ}$). Relaxing the requirements on energy resolution by a factor 2 (magenta curve) still allows for sensitivities above $3\sigma$ for exposures above 200 kt MW yr. For energy resolutions even better than our default assumption correspondingly smaller exposures are needed. We observe that the sensitivity saturates by resolutions better than a factor 2 compared to our default values.
@@ 5 $X_{T}$ test
The method to search for T violation used above is based on the presence of an $L$-odd component in the oscillation probability. As a variation of this, in our previous work [29] we have introduced a model-independent observable, called $X_{T}$ which is constructed from the difference of oscillation probabilities at two baselines. Here we briefly review this idea, specializing again to the standard three-flavour scenario and show numerical results.
@@
is negative, T violation must be present in the fundamental theory [29].
Figure 7: $X_{T}=P_{\nu_{\mu}\to\nu_{e}}(L_{\rm DUNE})-P_{\nu_{\mu}\to\nu_{e}}(L_{\rm T2HK})$ as a function of neutrino energy for neutrinos (left) and anti-neutrinos (right). For the oscillation parameters the best values from [9] have been used, assuming normal mass ordering. The orange and blue bands correspond to $X_{T}^{\rm even}$ implied by T-conservation, where the width of the band emerges from varying the lepton mixing angles within the current $3\sigma$ ranges. The solid-black (dashed-black) curves correspond to the case of maximal T violation with $\delta_{\rm CP}=90^{\circ}(270^{\circ})$. The purple-shaded band indicates the region where the conditions eqs. 5.5 and 5.6 are fulfilled simultaneously.
@@
Furthermore, we observe from fig. 7 that in the relevant energy range, $X_{T}$ for $\delta_{\rm CP}=90^{\circ}(270^{\circ})$ is negative (positive) for neutrinos, but positive (negative) for anti-neutrinos. Hence, using a neutrino beam, in the standard scenario this test can be applied for values of $\delta_{\rm CP}$ around $90^{\circ}$ [28, 29], where the T-violating value of $X_{T}$ has the opposite sign as for T conservation. Values around $270^{\circ}$ could be tested in principle with anti-neutrinos, although in practice the sensitivity is strongly reduced due to the large neutrino contamination in the “anti-neutrino” beam [29]. Therefore, we focus on the neutrino-beam mode for the numerical sensitivity analysis.
In our previous work [29] we developed a significance test based on the observable $X_{T}$. We applied this test without imposing any restriction on the coefficients $c_{2}$ and $c_{3}$, with the ambition of remaining as model-independent as possible. By marginalizing over $c_{2}$ and $c_{3}$, the right-hand side of eq. 5.3 can be made small and the significance of the test is given approximately by $|X_{T}^{\rm obs}|/\sigma_{X_{T}}$, with $\sigma_{X_{T}}$ denoting the uncertainty on $X_{T}^{\rm obs}$. In the present work we will take into account, that within the standard three-flavour framework the possible values of $c_{2}$ and $c_{3}$ are restricted by the allowed values of the oscillation parameters, implying that the right-hand side of eq. 5.3 will be constrained to sizeable non-zero values, increasing the significance to
@@
The finite energy resolution will lead to a smearing of the probability between bins. This can lead to the situation that the conditions eqs. 5.5 and 5.6 may no longer be fulfilled for energy-averaged quantities, when the chosen energy interval in eq. 5.8 refers to the true neutrino energy. In order to avoid this effect, we assume for the following calculation an energy resolution for DUNE a factor of 2 better than our default configuration. (This corresponds to the green curve in fig. 4.) See also [29] for a discussion of energy resolution effects. Finally, for sensitivity estimates, there is an ambiguity in whether we use the assumed true event numbers or the fitted event numbers for $S_{x}$ and $P_{x}$ to calculate $\sigma_{X_{T}}$; both options are frequently used in the literature. Below we will show results for both approaches.
@@ Appendix A Energy dependence of T and CP sensitivities
The left panel of fig. 9 shows the DUNE sensitivity to T violation (neutrino data only) for a wide range of energy bins. We see that for the chosen exposure, several bins provide sensitivity close to $2\sigma$, but the bin of $[0.80,0.92]$ GeV (shown in black) clearly sticks out with a sensitivity well above $2\sigma$. To gain further understanding of this result, we can consider an approximate expression for the transition probability in constant matter [54, 55, 56]:
$P_{\nu_{\mu}\to\nu_{e}}$ $\approx\sin^{2}2\theta_{13}\,\sin^{2}\theta_{23}\,\frac{\sin^{2}\Delta(1-A)}{(1-A)^{2}}+\alpha^{2}\sin^{2}2\theta_{12}\,\cos^{2}\theta_{23}\frac{\sin^{2}A\Delta}{A^{2}}$
$+\alpha\,\sin 2\theta_{13}\,\sin 2\theta_{12}\,\sin 2\theta_{23}\cos(\Delta+\delta_{\rm CP})\,\frac{\sin\Delta A}{A}\,\frac{\sin\Delta(1-A)}{1-A}\,$, (A.1)
@@
where $V$ is the effective matter potential [39]. Hence, the amplitude of the T-odd contribution, i.e., the coefficient in front of $\sin\delta_{\rm CP}$, is given by
$\mathcal{A}_{\rm odd}=-\alpha\,\sin 2\theta_{13}\,\sin 2\theta_{12}\,\sin 2\theta_{23}\sin\Delta\,\frac{\sin\Delta A}{A}\,\frac{\sin\Delta(1-A)}{1-A}\,.$ (A.3)

### v2

@@ 3 T violation from $L$ dependence
Further details on bin-wise and individual sensitivities are given in appendix A. Let us emphasize the crucial importance of the low-energy tail of the DUNE event spectrum for this analysis, covering the 2nd oscillation maximum. It is well known that the second oscillation maximum provides high sensitivity to $\delta_{\rm CP}$, see, for instance,[Rout:2020emr, Wildner:2015yaa, Qian:2013nhp, Coloma:2011pg, Ishitsuka:2005qi, Schwetz:2021cuj]. Here we argue, that this is a manifestation of T violation, as no neutrino/anti-neutrino comparison is involved but only the $L$ dependence of the oscillation probability (see further discussions in section 4 and appendix A).
Figure 1 illustrates also the dramatic impact of prior knowledge of the oscillation parameters on the sensitivity. For the green curves, no prior information on the mixing angles is assumed and we obtain a marginal sensitivity well below $2\sigma$. This situation is similar (though not identical) to the model-independent approach of our previous paper [Chatterjee:2024jzt]. [^1] Note that in [Chatterjee:2024jzt] we have assumed a DUNE exposure of 840 kt MW yr in order to boost the sensitivity, compared to the default exposure of 336 kt MW yr assumed here. In contrast, if a prior corresponding to the uncertainties from the current global fit on $\sin^{2}\theta_{ij}$ is imposed (blue curves in fig. 1), the sensitivity increases drastically, reaching more than $3\sigma$ for $\delta_{\rm CP}\approx 90^{\circ}$ for individual bins and more than $4\sigma$ for the combination. We note that at the time of the DUNE and T2HK experiments, an even better accuracy would be available from JUNO on $\theta_{12}$ and $\theta_{13}$ [JUNO:2022mxj] and from the DUNE/T2HK disappearance channels on $\theta_{23}$. Comparison of the blue and red curves in fig. 1 shows that a perfect knowledge of the mixing angles leads only to a slight improvement compared to present uncertainties.
@@
The left panel of fig. 3 shows that indeed the knowledge of all three mixing angles is important to reach such good sensitivities as displayed in fig. 1, especially on $\theta_{13}$ and $\theta_{23}$. Furthermore, an important effect comes also from the $\theta_{23}$ octant degeneracy. In our default analysis we assume a best fit value in the second octant, $\sin^{2}\theta_{23}=0.561$, and impose a prior on $\sin^{2}\theta_{23}$, which does determine the octant and therefore breaks the well-known octant degeneracy [Fogli:1996pv]. In the right panel of fig. 3 we study this issue in some detail. For the red curves we assume a prior on $\sin^{2}\theta$ (breaking the degeneracy), whereas for the blue curves the prior is imposed on $\sin^{2}2\theta$, which is symmetric with respect to the octant. By comparing the red and blue curves we see that the knowledge of the $\theta_{23}$ octant is rather important, in particular for the two combinations of true values ($\theta_{23}>45^{\circ}$, $\delta_{\rm CP}\simeq 90^{\circ}$) and ($\theta_{23}<45^{\circ}$, $\delta_{\rm CP}\simeq 270^{\circ}$). If the octant is known, the sensitivity is somewhat better for $\theta_{23}^{\rm true}<45^{\circ}$. While fig. 3 emphasizes the important to resolve the octant degeneracy, we note that the experiments considered here have good sensitivity to indeed lift this degeneracy, see e.g. [Ballett:2016daj, Chatterjee:2017irl, DUNE:2020jqi, Agarwalla:2024kti]. Therefore, we adopt the prior on $\sin^{2}\theta_{23}$ as our default assumption, assuming that the octant is known by the time the data becomes available.
@@
As the considered energy range is in the low-energy tail for DUNE, small event numbers are expected. Therefore, we study the sensitivity as a function of the DUNE exposure in fig. 4, together with the assumed energy resolution. We have checked that the results depend only mildly on both, the exposure and energy resolution in T2HK. For our default assumption on the energy resolution [Friedland:2018vry] as quoted in section 2.2 (blue curve), sensitivities above $4\sigma$ are obtained for exposures above 250 (400) kt MW yr in the neutrino running mode for a true $\delta_{\rm CP}=90^{\circ}$ ($270^{\circ}$). Relaxing the requirements on energy resolution by a factor 2 (magenta curve) still allows for sensitivities above $3\sigma$ for exposures above 200 kt MW yr. For energy resolutions even better than our default assumption correspondingly smaller exposures are needed. We observe that the sensitivity saturates by resolutions better than a factor 2 compared to our default values.
@@ 5 $X_{T}$ test
The method to search for T violation used above is based on the presence of an $L$-odd component in the oscillation probability. As a variation of this, in our previous work [Chatterjee:2024jzt] we have introduced a model-independent observable, called $X_{T}$ which is constructed from the difference of oscillation probabilities at two baselines. Here we briefly review this idea, specializing again to the standard three-flavour scenario and show numerical results.
@@
is negative, T violation must be present in the fundamental theory [Chatterjee:2024jzt].
Figure 7: $X_{T}=P_{\nu_{\mu}\to\nu_{e}}(L_{\rm DUNE})-P_{\nu_{\mu}\to\nu_{e}}(L_{\rm T2HK})$ as a function of neutrino energy for neutrinos (left) and anti-neutrinos (right). For the oscillation parameters the best values from [Esteban:2024eli] have been used, assuming normal mass ordering. The orange and blue bands correspond to $X_{T}^{\rm even}$ implied by T-conservation, where the width of the band emerges from varying the lepton mixing angles within the current $3\sigma$ ranges. The solid-black (dashed-black) curves correspond to the case of maximal T violation with $\delta_{\rm CP}=90^{\circ}(270^{\circ})$. The purple-shaded band indicates the region where the conditions eqs. 5.5 and 5.6 are fulfilled simultaneously.
@@
Furthermore, we observe from fig. 7 that in the relevant energy range, $X_{T}$ for $\delta_{\rm CP}=90^{\circ}(270^{\circ})$ is negative (positive) for neutrinos, but positive (negative) for anti-neutrinos. Hence, using a neutrino beam, in the standard scenario this test can be applied for values of $\delta_{\rm CP}$ around $90^{\circ}$ [Schwetz:2021cuj, Chatterjee:2024jzt], where the T-violating value of $X_{T}$ has the opposite sign as for T conservation. Values around $270^{\circ}$ could be tested in principle with anti-neutrinos, although in practice the sensitivity is strongly reduced due to the large neutrino contamination in the “anti-neutrino” beam [Chatterjee:2024jzt]. Therefore, we focus on the neutrino-beam mode for the numerical sensitivity analysis.
In our previous work [Chatterjee:2024jzt] we developed a significance test based on the observable $X_{T}$. We applied this test without imposing any restriction on the coefficients $c_{2}$ and $c_{3}$, with the ambition of remaining as model-independent as possible. By marginalizing over $c_{2}$ and $c_{3}$, the right-hand side of eq. 5.3 can be made small and the significance of the test is given approximately by $|X_{T}^{\rm obs}|/\sigma_{X_{T}}$, with $\sigma_{X_{T}}$ denoting the uncertainty on $X_{T}^{\rm obs}$. In the present work we will take into account, that within the standard three-flavour framework the possible values of $c_{2}$ and $c_{3}$ are restricted by the allowed values of the oscillation parameters, implying that the right-hand side of eq. 5.3 will be constrained to sizeable non-zero values, increasing the significance to
@@
The finite energy resolution will lead to a smearing of the probability between bins. This can lead to the situation that the conditions eqs. 5.5 and 5.6 may no longer be fulfilled for energy-averaged quantities, when the chosen energy interval in eq. 5.8 refers to the true neutrino energy. In order to avoid this effect, we assume for the following calculation an energy resolution for DUNE a factor of 2 better than our default configuration. (This corresponds to the green curve in fig. 4.) See also [Chatterjee:2024jzt] for a discussion of energy resolution effects. Finally, for sensitivity estimates, there is an ambiguity in whether we use the assumed true event numbers or the fitted event numbers for $S_{x}$ and $P_{x}$ to calculate $\sigma_{X_{T}}$; both options are frequently used in the literature. Below we will show results for both approaches.
@@ Appendix A Energy dependence of T and CP sensitivities
The left panel of fig. 9 shows the DUNE sensitivity to T violation (neutrino data only) for a wide range of energy bins. We see that for the chosen exposure, several bins provide sensitivity close to $2\sigma$, but the bin of $[0.80,0.92]$ GeV (shown in black) clearly sticks out with a sensitivity well above $2\sigma$. To gain further understanding of this result, we can consider an approximate expression for the transition probability in constant matter [Cervera:2000kp, Freund:2001pn, Akhmedov:2004ny]:
$P_{\nu_{\mu}\to\nu_{e}}$ $\approx\sin^{2}2\theta_{13}\,\sin^{2}\theta_{23}\,\frac{\sin^{2}\Delta(1-A)}{(1-A)^{2}}+\alpha^{2}\sin^{2}2\theta_{12}\,\cos^{2}\theta_{23}\cos^{2}\theta_{13}\frac{\sin^{2}A\Delta}{A^{2}}$
$+\alpha\,\cos\theta_{13}\sin 2\theta_{13}\,\sin 2\theta_{12}\,\sin 2\theta_{23}\cos(\Delta+\delta_{\rm CP})\,\frac{\sin\Delta A}{A}\,\frac{\sin\Delta(1-A)}{1-A}\,$, (A.1)
@@
where $V$ is the effective matter potential [Wolfenstein:1977ue]. Hence, the amplitude of the T-odd contribution, i.e., the coefficient in front of $\sin\delta_{\rm CP}$, is given by
$\mathcal{A}_{\rm odd}=-\alpha\,\cos\theta_{13}\sin 2\theta_{13}\,\sin 2\theta_{12}\,\sin 2\theta_{23}\sin\Delta\,\frac{\sin\Delta A}{A}\,\frac{\sin\Delta(1-A)}{1-A}\,.$ (A.3)
@@
Appendix B Estimating corrections for variable matter density
All results in this paper assume a constant, line-averaged matter density of $\bar{\rho}=2.84\,\rm g/cm^{3}$ for both T2HK and DUNE. While this matches the DUNE baseline value, it differs slightly from the T2HK average of $2.6\,\rm g/cm^{3}$. In reality, both experiments feature non-constant, asymmetric density profiles [Hagiwara:2011kw, Roe:2017zdw]. This appendix details the impact of matter density on CP and T violation sensitivities by comparing two scenarios: (I) constant but different averaged densities for each experiment, and (II) the real, variable density profiles along each baseline.
Figure 12: Relative corrections to the $\nu_{\mu}\rightarrow\nu_{e}$ appearance probability ($P_{\mu e}$) for T2HK (left) and DUNE (right) as a function of the CP-phase $\delta_{\rm CP}$. Results are shown for two energies, $E_{\nu}=0.74$ and $0.86$ GeV. Solid lines represent the realistic non-constant density profiles [Hagiwara:2011kw, Roe:2017zdw], while dashed lines show the effect of using different constant average densities.
Figure 12 shows the relative error in $P_{\mu e}$ induced by these density choices. For T2HK (left panel, dashed lines), using the DUNE-averaged density ($\bar{\rho}_{\rm DUNE}=2.84\,\rm g/cm^{3}$) instead of its own ($\bar{\rho}_{\rm T2HK}=2.6\,\rm g/cm^{3}$) results in a negligible error of less than $1\%$ across all $\delta_{\rm CP}$ values. This corresponds to a change of less than one event per hundred in a typical bin, justifying its neglect in our sensitivity analysis. This case is omitted for DUNE (right panel) as the relative error is zero by definition when $\bar{\rho}_{\rm DUNE}=\bar{\rho}$.
We now consider the realistic profiles from [Hagiwara:2011kw, Roe:2017zdw] (Case II, solid lines). Despite T2HK’s step-like asymmetric potential and DUNE’s close to symmetric profile, the resulting corrections remain remarkably small. The error consistently stays below $1\%$, with a single exception: for DUNE at $E_{\nu}=0.86$ GeV and $\delta_{\rm CP}\approx 90^{\circ}$, it reaches $\sim 1.5\%$. These deviations are negligible in terms of expected event counts.
In conclusion, the use of a unified constant density of $2.84\,\rm g/cm^{3}$ is well-justified for the comparative analysis of T2HK and DUNE presented in this work, as the introduced errors are completely insignificant compared to statistical uncertainties, see e.g., fig. 2 for the typical size of error bars in the relevant bins.